OpenStudy (anonymous):

Can I please get help with these data questions? 1) Alice's average mark for five exams is 67%. What minimum mark should Alice obtain in her sixth exam if her average mark for the six exams is to be 70%? 2) For the following set of scores, write another score so that the range is equal to the mode: 19, 21, 23, 23, 23, 27, 34

OpenStudy (anonymous):

1) In 5 exams her average is 67.So her total marks is 67*5=335 for all of them.. After the 6th exam she needs to get 70 % average so her marks should be 6*70=420. So she needs to get 420-335=85% marks in her 6th exam 2)19,21,23,23,23,27,34 mode=23=range since range=23, difference between highest and lowest marks is 23. So the 8th score can be either the lowest or the highest if it is highest then 19 is lowest so score=19+23=42 if it is lowest then 34 is highest so 34-23=11 is the score so u have 2 possible answers

OpenStudy (anonymous):

Thanks viktor.ray! All I wanted was a hint but you did it all. Thanks!